OpenStudy (anonymous):

differentiate xylog(x+y)=1

OpenStudy (psymon):

Is this log base 10 or base e?

OpenStudy (goformit100):

it it N.C.E.R.T. Question ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

@Psymon its not mentioned

OpenStudy (psymon):

I see. Well, we can try it with base e or base 10, which do you think is more likely?

OpenStudy (anonymous):

base 10

OpenStudy (psymon):

Alright then. So log base 10 derivative is: \[\frac{ 1 }{ (\ln10)u }*\frac{ du }{ dx}\] So this is going to be a product rule of 3 terms \[f'(x)g(x)h(x) + f(x)g'(x)h(x) + f(x)g(x)h'(x)\] Does it say if we have to differentiate with respect to x or y by chance?

OpenStudy (anonymous):

w.r.t x

OpenStudy (psymon):

Alright, just making sure :3 Although it looks like genius12 has got theproblem cooking up, lol. Let's see what hecomes up with xD

OpenStudy (anonymous):

actually this is a proof dy/dx=-y(x^2y+x+y)/x(xy^2+x+y)

OpenStudy (psymon):

Ah. Either way, I'm sure genius has got it xD No need for me to interrupt his work.

OpenStudy (anonymous):

lol:D

OpenStudy (anonymous):

no one got the answer???

OpenStudy (psymon):

o.o no-one posted an answer O.o Um, I could try....as long as Im not interrupting someone else.

OpenStudy (anonymous):

pls try

OpenStudy (anonymous):

how will we get dy if we differentiate w.r.t x?

OpenStudy (psymon):

Alrighty. So the first portion of the product rule would be this then: \[(1)(y)\log(x+y)\] Next we would have: \[(x)(1)\frac{ dy }{ dx }\log(x+y)\] and the 3rd term: \[(x)(y)\frac{ 1 }{ (\ln10)(x+y) }*(1+\frac{ dy }{ dx })\] So now we need to separate our dy/dx terms from the other terms.

OpenStudy (psymon):

Since we have a dy/dx inside of a parenthesis in that 3rd term, I have to multiply that through: \[\frac{ xy }{ (\ln10)(x+y) }+\frac{ xy }{ (\ln10)(x+y) }*\frac{ dy }{ dx }\]

OpenStudy (psymon):

We have 2 terms with dy/dx and 2 terms without it, so I'll separate those on opposite sides of the equal sign \[\frac{ dy }{ dx }( xlog(x+y)+\frac{ xy }{ (\ln10)(x+y) }) = -(ylog(x+y) + \frac{ xy }{ (\ln10)(x+y) }\] Now I need to divide both sides by what is with dy/dx

OpenStudy (psymon):

\[-(ylog(x+y) + \frac{ xy }{ (\ln10)(x+y) })*(\frac{ 1 }{ (xlog(x+y)) }+\frac{ (\ln10)(x+y) }{ xy })\] Now let me simplify this on paper, haha.

OpenStudy (psymon):

Yeah, looking at your proof now, this must be log base e O.o

OpenStudy (anonymous):

actually this is a proof dy/dx=-y(x^2y+x+y)/x(xy^2+x+y)

OpenStudy (psymon):

Yeah, I noticed it now, which makes me think this would be log base e and not log base 10.

OpenStudy (anonymous):

ssry my mistake

OpenStudy (anonymous):

u were tying smthing wht happened

OpenStudy (psymon):

I guess I'm wondering where all the ln's went. I'm not sure how to get them to all disappear in your proof like that. Maybe I'm not sure how to do it then. I mean, I solve for (dy/dx), but it's nothing like that \[\frac{ dy }{ dx } =\frac{ -y[(x+y)\ln(x+y)+x] }{ x[(x+y)\ln(x+y)+y] }\] I must flat out be missing something, lol x_x

OpenStudy (anonymous):

i know even i dont know where log vanished i tried this sum 5 times still i didn't get

OpenStudy (psymon):

It won't disappear for log or ln. My calculator gave me a middle finger when I tried with that. Wish i could say I knew where that came from O.o

OpenStudy (anonymous):

wht i should do now

OpenStudy (anonymous):

this question is there in RD sharma but unsolved

hartnn (hartnn):

did anyone try substituting x+y= u ?

OpenStudy (psymon):

Would that actually get rid of the log when you finish, though? I would think some of the terms would still have the log even if you did that substitution.

hartnn (hartnn):

yep

hartnn (hartnn):

i tried and i got rid of log

hartnn (hartnn):

as of now i am here, simplification is remaining 1+ dy/dx = (x+y) d/dx (1/(xy))

OpenStudy (psymon):

I see. So would you have to make y = u-x as well then?

hartnn (hartnn):

what i did is x+y= u (just for convenience, we can do it without doing this too) 1+ dy/dx = du yep, y=u-x x(u-x) log u = 1 u= e^ (1/ [x u-x]) now differentiate both sides w.r.t x

hartnn (hartnn):

***u= e^ (1/ [x (u-x)])

hartnn (hartnn):

shalini, are you following so far ?

OpenStudy (anonymous):

not exactly

hartnn (hartnn):

ok, which step you have doubt in ?

OpenStudy (anonymous):

first u complete

OpenStudy (anonymous):

i understood

hartnn (hartnn):

eh, complete ? i want you to learn, by interaction....giving you entire solution at once would not help much....try to get each and every step

hartnn (hartnn):

and i gave you a headd start. i would be glad if you can think of next steps...

OpenStudy (anonymous):

i understood till where u hv done

hartnn (hartnn):

so, can you differentiate w.r.t x that for me ?? u= e^ (1/ [x (u-x)]) du/dx =.. ?

OpenStudy (anonymous):

sir,i m very bad in differentiation

OpenStudy (anonymous):

i will try

hartnn (hartnn):

what id d/dx of e^x ?

OpenStudy (anonymous):

its e^x

hartnn (hartnn):

good, now lets see whether you know chain rule, what is d/dx of e^x^2 ?

OpenStudy (anonymous):

2xe^x^2

hartnn (hartnn):

oh wow, you know chain rule :) so you can surely and easily differentiate this. d/dx (e^anything) = e^anything d/dx (anything) u= e^ (1/ [x (u-x)]) du/dx =.. ?

hartnn (hartnn):

just one next step, not entire thing..

OpenStudy (anonymous):

w8

OpenStudy (psymon):

Is this supposed to come out as complicated as I'm getting it? I mean, I've done the derivative and isolated du/dx, but it's looking a bit insane O.o

hartnn (hartnn):

thats why i asked to give me just one next step....you need not derivate the remaining part or isolate "du/dx"

OpenStudy (psymon):

Well, I didn't want to post steps for him, I was just trying to do it on my own, sorry.

OpenStudy (anonymous):

i think du/dx= e^1/x(u-x)*-[u-2x]/(x(u-x))^2

hartnn (hartnn):

don't be sorry, we are learning....you can message me the next step, and i will tell you what u can do further

hartnn (hartnn):

ehh....see this became complicated...but ok, what i was expecting is this, du/dx = e^ (1/ [x (u-x)]) d/dx [e^ (1/ [x (u-x)])] but we have e^ (1/ [x (u-x)]) = u !!! so, du/dx = u d/dx (e^ (1/ [x (u-x)])) got this ?

hartnn (hartnn):

***du/dx =u d/dx(1/ [x(u-x)])

hartnn (hartnn):

now we don't need u, we can bring back our 'y' 1+dy/dx = (x+y) d/dx (1/xy) now this isn't very complicated

hartnn (hartnn):

this was important step i guess, so any of you have any doubts, please get it cleared

OpenStudy (psymon):

Still staring at it, haha.

OpenStudy (anonymous):

no continue

OpenStudy (psymon):

Alright, I guess I see what ya did.

hartnn (hartnn):

let me write all steps together, i made some typos earlier u= e^ (1/ [x (u-x)]) du/dx = e^ (1/ [x (u-x)]) d/dx [(1/ [x (u-x)])] but u= e^ (1/ [x (u-x)]) so, du/dx = u d/dx[1/x(u-x)] bringing back 'y' , {u=x+y so, du/dx = 1+dy/dx and u-x=y} 1+dy/dx = (x+y) d/dx (1/xy)

hartnn (hartnn):

shalini, i guess you know both chain and product rule, so can you try this for me, separately ?? d/dx (1/xy)

OpenStudy (psymon):

No need to differentiate the (x+y)?

hartnn (hartnn):

nopes...because of chain rule we had to diff. 1/(xy)

OpenStudy (psymon):

So the (x+y) will just be there to multiply with the result of d/dx(1/xy).

hartnn (hartnn):

correct

OpenStudy (anonymous):

-xdy/dx+y/x^2y^2

hartnn (hartnn):

did you mean \(\huge -\dfrac{xy'+y}{x^2y^2}\) ?? where y' = dy/dx , just for convenience

OpenStudy (anonymous):

yes y its not correct

hartnn (hartnn):

so now you have \(\huge 1+y'=\huge -(x+y)(\dfrac{xy'+y}{x^2y^2})\) try to isolate y', this work is purely algebraic

hartnn (hartnn):

oh and i forgot to ask, any doubts till that step ?

OpenStudy (psymon):

Alright, so since we got to that point I'll now ask my question, lol. So I would have: \[-(\frac{ 1 }{ x ^{2}y } + \frac{ xy' }{ (xy)^{2} })\] Now because the x is being multiplied by y', am I not allowed to cancel it out with one of the x's in the denominator or can I?

hartnn (hartnn):

you can, surely, xy' / x^2y^2 = y'/ xy^2

OpenStudy (psymon):

Alright, making sure :3

OpenStudy (anonymous):

no doubt

hartnn (hartnn):

good, so i think distributing x+y over that big bracket, like psymon did , will be a good step to start simplifying...

hartnn (hartnn):

remember our aim is to isolate y'

hartnn (hartnn):

then the next good step, to get rid of the denominator, will be to multiply both sides by x^2 y^2

OpenStudy (anonymous):

i didn't get Psymon's step

hartnn (hartnn):

then try it by yourself ?

OpenStudy (anonymous):

arre why r u getting angry

hartnn (hartnn):

angry ? lol, not at all...if i don't put smileys doesn't mean i am getting angry i am just encouraging you to first try it by yourself, if you don't get it, we are here to help you :)

OpenStudy (psymon):

That turned out pretty awesome x_x And my step was instinctive to just try and start to isolate y'. After the quotient rule, I split the result into two separate fractions. That was all i did : )

OpenStudy (anonymous):

from where psymon got −(1x2y+xy′(xy)2)

hartnn (hartnn):

how will you simplify (x+y)*z ?

OpenStudy (anonymous):

xz+yz

OpenStudy (psymon):

I'll show ya what I did. \[-(\frac{ xy'+y }{ x ^{2}y ^{2} })\] So this was the result of our quotient rule. For every term in the numerator, I can make a new fraction, using that term over the whole denominator. So I split it into: \[-(\frac{ xy' }{ x ^{2}y ^{2} }+\frac{ y }{ x ^{2}y ^{2} })\] Even thoughthis is what I did, don't let it confuse you I just insticitvely broke it apart, it is probably more efficient to do what hartnn said and just multiply out x^2y^2 from the beginning xD

OpenStudy (anonymous):

k

hartnn (hartnn):

so, (x+y) (xy′+y/x^2y^2) will be just =x (xy′+y/(x^2y^2)) + y (xy′+y/(x^2y^2)) right ? yep,all have different ways to simply things, i will surely multiply x^2y^2 right now to get rid of denominator

OpenStudy (psymon):

I agree, just first was my first thought, doesn't mean it was most efficient :P